OpenStudy (anonymous):

another one for ya...solve. 3x+6 (10-9x) = 3 (20-17x)

OpenStudy (anonymous):

3x + 60 - 54x = 60 - 51x 60 - 51x = 60 - 51x 60 - 60 = -51x + 51x 0 = 0 so x = 0

OpenStudy (anonymous):

0 = 0 ---> All x allowed

OpenStudy (anonymous):

Scallmann is correct....... I overlooked that aspect...